OpenStudy (anonymous):

The sales at a particular bookstore grew from $2006 million in 2000 to $3803 million in 2005. Find an exponential function to model the sales as a function of years since 2000. Give your answer using the form B = Boat.

OpenStudy (anonymous):

\[B=B_oa^t\]

OpenStudy (anonymous):

yea use the formula

OpenStudy (anonymous):

\[B_o=2006\]

OpenStudy (anonymous):

but how do i find a?

OpenStudy (helder_edwin):

t=0 corresponds to 2000 then t=? for 2005

OpenStudy (anonymous):

5

OpenStudy (helder_edwin):

yes then u have \[ \large 3803=B_0a^5=2006\cdot a^5 \]

OpenStudy (anonymous):

3806/2006=a^5

OpenStudy (anonymous):

how do i find a after this?

OpenStudy (anonymous):

log?

OpenStudy (helder_edwin):

yes

OpenStudy (anonymous):

idk how

OpenStudy (anonymous):

\[\log(3806/2006)=\log(a)*5\]

OpenStudy (anonymous):

can I cancel the logs?

OpenStudy (helder_edwin):

no. u just put them. divide by 5

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

\[(\log(3806/2006))/(\log(a))=5\]

OpenStudy (helder_edwin):

no

OpenStudy (anonymous):

I dont know what to do

OpenStudy (helder_edwin):

\[ \large \log(3806/2006)=5\log a \] \[ \large \frac{\log(3806/2006)}{5}=\log a \]

OpenStudy (helder_edwin):

then \[ \large \log a=\log(3806/2006)^{1/5} \]

OpenStudy (helder_edwin):

got it??

OpenStudy (anonymous):

ummm i think so

OpenStudy (anonymous):

but whats a?

OpenStudy (anonymous):

i need to find a

OpenStudy (helder_edwin):

sorry. now u can cancel the logs

OpenStudy (helder_edwin):

\[ \large a=(3806/2006)^{1/5} \]

OpenStudy (helder_edwin):

got it?? (Megan Fox)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thats not the right answer :/

OpenStudy (anonymous):

i need to find the right a

OpenStudy (anonymous):

"a"

OpenStudy (anonymous):

what do you get for a? i got 1.136

OpenStudy (anonymous):

that answer for "a" is wrong and i dont know why

OpenStudy (anonymous):

i need an exact answer

OpenStudy (anonymous):

 but its wrong